OpenStudy (anonymous):

heres the question it deals with complex number yet I dont know exactly whats its asking. Let H(\omega) be a complex-valued function of the real variable \omega. Find |H(\omega)| and argH(\omega) equations below

OpenStudy (anonymous):

\[H(\omega) = \frac{ 1 }{ (1+i\omega)^{10}}\]

OpenStudy (cwrw238):

i think you need to convert it to polar form r(cosx +isinx) where r = modulus and x is the argument

OpenStudy (anonymous):

ok so \[\left| H(\omega) \right|\] is r?

OpenStudy (anonymous):

yes, but i'm not sure if you have to work it out

OpenStudy (anonymous):

i'm currently working out the argument, and i think i've got it too

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

there's several rules for arguments \[\large \arg(z^n)=n~\arg(z)\] \[\large \arg(\frac{ z_1 }{ z_2 })=\arg(z_1)-\arg(z_2)\] \[\large \arg(x+yi)=\tan^{-1}(\frac{ y }{ x })\]

OpenStudy (anonymous):

if i'm not mistaken that would mean we have arg(1)-10 arg(1+i\(\omega\))

OpenStudy (anonymous):

which would be 0-tan\(^{-1}(\omega)\) or simply -tan\(^{-1}(\omega)\)

OpenStudy (anonymous):

ok Im trying to follow.

OpenStudy (anonymous):

actually i forgot the 10 in the last one, so \(-10tan^{-1}(\omega)\)

OpenStudy (anonymous):

Let \(z=1+i\omega\), then \[\begin{cases} r\cos\theta=1\\ r\sin\theta=\omega \end{cases}~~\implies~~r^2\cos^2\theta+r^2\sin^2\theta=1+\omega^2=r^2\] so \(r=\sqrt{1+\omega^2}\). \[\theta=\tan^{-1}\frac{y}{x}=\tan^{-1}\omega\] So you have \[\begin{align*}H(\omega)&=\frac{1}{(1+i\omega)^{10}}\\\\ &=\frac{1}{\left(\sqrt{1+\omega^2}\exp\left(i\tan^{-1}\omega\right)\right)^{10}}\\\\ &=\frac{1}{\left(1+\omega^2\right)^5\exp\left(10i\tan^{-1}\omega\right)}&\text{ by DeMoivre's theorem} \end{align*}\] Try taking the modulus now.

OpenStudy (anonymous):

thanks guys!